A 800-N bear walks out on a beam in an attempt to retrieve some “goodies” hanging at the end (Pictured below). The beam is uniform, weighs 250 N, and is 6.00 m long; the goodies weigh 70.0 N. If the wire can withstand a maximum tension of 900 N, what is the maximum distance the bear can walk before the wire breaks?
Answer #1

mbear x + mbeam (L/2) + mbasket L = T sin(theta) L

(800 x X) + (250 x 0.5 x 6) + (70 x 6) = 900 x sin(60) x 6

800 X + 750 + 420 = 4676.54

X = 4.38 m

Comment in case any doubt please rate my answer....